OpenStudy (anonymous):

limit as x approaches 0 (x^2-sin^2(3x)) / 2x^2

OpenStudy (taylorterminate):

lim (x-->0) sin²(3x)/x² = lim (x-->0) [sin(3x)/x]² = [lim (x-->0) sin(3x)/x]² = [3 * lim (x-->0) sin(3x)/(3x)]² = (3 * 1)² = 9.

OpenStudy (anonymous):

no, the answer is -4, I just don't know how to find it. Can someone help me?

OpenStudy (anonymous):

\[\begin{align*}\frac{x^2-\sin^23x}{2x^2}&=\frac{1}{2}-\frac{\sin^23x}{2x^2}\\ &=\frac{1}{2}-\frac{9\sin^23x}{2(9x^2)}\\ &=\frac{1}{2}-\frac{9}{2}\left(\frac{\sin 3x}{3x}\right)^2 \end{align*}\] As \(x\to0\), the \(\dfrac{\sin3x}{3x}\) term approaches 1, leaving you with \(\dfrac{1}{2}-\dfrac{9}{2}=-4\).

OpenStudy (anonymous):

thank you!!